OpenStudy (anonymous):

Jenny, who rides a moped, takes 2 hours less to travel 60 miles than Maureen takes to travel 50 miles on her bicycle. Jenny travels 10 miles per hour faster than Maureen. (Hint: speed = distance ÷ time.)

OpenStudy (anonymous):

I just don't know how to plug in what I have.

OpenStudy (anonymous):

60x - 50(x-2) = 10x(x-2) ----------------------- 10x + 100 = 10x^2 -20x 10x^2 -30x - 100 = 0 x^2 - 3x - 10 = 0 (x-5)(x+2) = 0 Positive solution:: x = 5 --------------------- Maureen DATA: time = x = 5 hrs ; rate = 50/5 = 10 mph Jenny DATA: time = x-2 = 3 hrs ; rate = 60/3 = 20 mph ===========================

OpenStudy (anonymous):

Maureen's speed = 50 / Jenny's time + 2 hrs. Jenny's speed = Maureen's speed + 10 miles/hour Jenny's speed = Maureen's time - 2 hours / 60 miles. I think that's what you're looking for.

OpenStudy (anonymous):

hope it help's .....

OpenStudy (anonymous):

Jenny, who rides a moped, takes 2 hours less to travel 60 miles than Maureen takes to travel 50 miles on her bicycle. Jenny travels 10 miles per hour faster than Maureen. (Hint: speed = distance ÷ time.) ------- Jenny DATA: distance = 60 miles ; time = x-2 hrs ; rate = d/t = 60/(x-2) mph ------- Maureen DATA: distance = 50 miles ; time = x hrs ; rate = 50/x mph --------------------- Equation: rate - rate = 10 mph 60/(x-2) - 50/x = 10 mph ----- 60x - 50(x-2) = 10x(x-2) ----------------------- 10x + 100 = 10x^2 -20x 10x^2 -30x - 100 = 0 x^2 - 3x - 10 = 0 (x-5)(x+2) = 0 Positive solution:: x = 5 --------------------- Maureen DATA: time = x = 5 hrs ; rate = 50/5 = 10 mph Jenny DATA: time = x-2 = 3 hrs ; rate = 60/3 = 20 mph ===========================

OpenStudy (anonymous):

@Xd1521 can you explain how you got from the question to your solution, because it sounds like what the person is having problems with is plugging things into the equation, and I'm not sure how you did it there, that might help them out.

OpenStudy (anonymous):

where i got from u mean? @PsychoTink

OpenStudy (anonymous):

yea that works too.("

OpenStudy (anonymous):

does that help at all?

OpenStudy (anonymous):

I can google. I see you very clearly got the answer from http://www.algebra.com/algebra/homework/word/travel/Travel_Word_Problems.faq.question.899068.html But that doesn't help teach the person who asked this question HOW to solve the problem theirself, where the numbers came from and why.

OpenStudy (anonymous):

i know i didnt know the ansewer thts y i googled

OpenStudy (anonymous):

easy look their and u can look it up easly .

OpenStudy (anonymous):

google is a thing for learn from ...

OpenStudy (anonymous):

but just got look for what your look for .

OpenStudy (anonymous):

@Icecube902 @YOUNGSUPERMAN

OpenStudy (youngsuperman):

yes

OpenStudy (e.mccormick):

Two things, "google is a thing for learn from ..." is an incorrect sentence and has no meaning. Second, while you can use Google to find an answer, that does not mean you are learning. Getting an answer and learning are separate things. If you learn how to set up and do the problem then you can easily get answers. Also, being able to translate these types of word problems helps you learn how to do real life applications of math in an assortment of fields. So it is far better to learn the material than just to find an answer.